### v3

# Roadside Monocular 3D Detection Prompted by 2D Detection

Abstract
Roadside monocular 3D detection requires detecting objects of predefined classes in an RGB frame and predicting their 3D attributes, such as bird’s-eye-view (BEV) locations. It has broad applications in traffic control, vehicle-vehicle communication, and vehicle-infrastructure cooperative perception. To address this task, we introduce Promptable 3D Detector (Pro3D), a novel detector design that leverages 2D detections as prompts. We build our Pro3D upon two key insights. First, compared to a typical 3D detector, a 2D detector is “easier” to train due to fewer loss terms and performs significantly better at localizing objects w.r.t 2D metrics. Second, once 2D detections precisely locate objects in the image, a 3D detector can focus on lifting these detections into 3D BEV, especially when fixed camera pose or scene geometry provide an informative prior. To encode and incorporate 2D detections, we explore three methods: (a) concatenating features from both 2D and 3D detectors, (b) attentively fusing 2D and 3D detector features, and (c) encoding properties of predicted 2D bounding boxes {$x$, $y$, width, height, label} and attentively fusing them with the 3D detector feature. Interestingly, the third method significantly outperforms the others, underscoring the effectiveness of 2D detections as prompts that offer precise object targets and allow the 3D detector to focus on lifting them into 3D. Pro3D is adaptable for use with a wide range of 2D and 3D detectors with minimal modifications. Comprehensive experiments demonstrate that our Pro3D significantly enhances existing methods, achieving state-of-the-art results on two contemporary benchmarks.
Figure 1: A summary of our work. (a) To approach roadside monocular 3D detection, we introduce a novel detector design Promptable 3D detector (Pro3D) that exploits 2D detections to prompt the 3D detector for 3D prediction. Pro3D can exploit any 2D detector and 3D detector with minimal modificaitons. (b) The Pro3D design is motivated by the observation that a simple 2D detector DINO [62] significantly outperforms state-of-the-art roadside monocular 3D detectors (e.g., BEVHeight [56] and BEVSpread [48]) w.r.t 2D metrics (on the DAIR-V2X-I dataset). This implies that 2D detection is an “easier” task than monocular 3D detection – with a trained 2D detector, training 3D detector can be “simplified” by learning to lift 2D detections to the 3D space. Moreover, as roadside camera pose is fixed, we derive a backbround image as scene prior, incorporating which remarkably boosts 3D detection performance. (c) As a summary of results, our Pro3D significantly outperforms prior works on the DAIR-V2X-I benchmark. Refer to Table 1 and 2 for comprehensive results.
1 Introduction
Roadside monocular 3D detection is a practical yet challenging task that requires, over a single RGB frame, detecting objects of predefined classes (e.g., vehicle and pedestrian) and predicting their 3D properties such as the locations, shape, and orientations in bird’s-eye-view (BEV) [56]. It has broad applications [61, 59] in vehicle-vehicle communication [9, 17], vehicle-infrastructure cooperative perception [1], and intelligent traffic control [33, 47].
Status quo. Roadside monocular 3D detection is an ill-posed problem as it requires inferring 3D information from 2D cues. Fortunately, camera pose is fixed on roadside infrastructure, making it possible to learn from data to infer 3D properties from a 2D image. To foster the research of roadside monocular 3D detection, the community has contributed large-scale benchmark datasets, such as DAIR-V2X-I [61] and Rope3D [59]. Existing methods train sophisticated deep neural networks in an end-to-end fashion [19, 18, 56, 48]. For monocular 3D detection, a notorious difficulty is precisely estimating depth for objects [52, 38, 12]. To mitigate this difficulty, prior works propose to convert depth estimation to object height estimation and exploits camera parameters to infer object depth [56], or consider scene geometry to carefully pool features for better 3D prediction [48].
Motivation. We observe that the state-of-the-art roadside 3D detectors BEVHeight [56, 58] and BEVSpead [48] significantly underperform 2D detectors (e.g., DINO [62]) w.r.t 2D detection metrics, as shown by the precision-recall curves on the DAIR-V2X-I dataset [61] in Fig. 1b. Here, we train the 2D detector DINO [62] on this dataset, using the generated 2D bounding boxes by projecting 3D cuboid annotations onto the image plane. To evaluate BEVHeight w.r.t 2D detection metrics, we project its 3D detections onto the image plane as its 2D detections. This comparison has several important implications: (1) 2D detection is an “easier” task than monocular 3D detection, and (2) the fixed camera pose or scene geometry provides an informative prior that greatly simplifies 2D detection. The results motivate us to exploit 2D detections and scene priors to facilitate 3D detector training, leading to our novel detector design (Fig. 1a), Promptable 3D Detector (Pro3D).
Technical insights. Training a 3D detector involves optimizing multiple objectives, such as 2D box coordinate regression, depth estimation, orientation regression, and object classification. Intuitively, a compound holistic objective that optimizes them all together might cause the missing detection of some hard objects as an attempt to handle the difficulty in regressing coordinates, depth and orientation. In contrast, training a 2D detector is simpler that incorporates fewer loss terms, i.e., only a 2D box coordinate regression loss and a classification loss. 2D detection is also well-studied in the literature that develops a multitude of useful techniques. These help a 2D detector outperform 3D detectors on 2D metrics, even when both are trained on the same data. This observation leads to our core insight that exploits a 2D detector to facilitate 3D detector training (Fig. 1a). Intuitively, 2D detections pinpoint the target objects in the image and the 3D detector can be simply thought of as learning to lift them to 3D BEV. From a contemporary view, our approach uses a 2D detector to prompt 3D detector, hence we name our framework Promptable 3D Detector (Pro3D). Moreover, as the camera is fixed in roadside perception, we derive a “blank” scene background as a scene prior from training frames (Fig. 4). We incorporate the scene prior as a part of input to the 3D detector. Our experiments validate the usefulness of exploiting the scene prior, and importantly, the design of Pro3D for roadside monocular 3D detection.
Contributions. We make three contributions.
  1. 1.
    We introduce a novel detector design Pro3D, which serves as a general framework for roadside monocular 3D detection (Fig. 1a). Pro3D is adaptable with any 2D detectors and roaside 3D detectors: it produces 2D detections, then uses them as prompts to help a 3D detector make better 3D detections. We develop multiple methods to exploit the 2D detector and derive the simple yet effective approach – attentively fusing 2D detections.
  2. 2.
    To better solve roadside monocular 3D detection, we present an embarrassingly simple method to obtain a scene prior: detecting and removing objects in frames, and averaging the pixels across frames towards a “blank” scene background as the scene prior. Incorporating the scene prior greatly enhances 3D detection.
  3. 3.
    We evaluate our Pro3D with different 2D detectors and 3D detectors through extensive experiments and ablation studies, showing that Pro3D significantly improves over existing roadside 3D detectors and achieves the state-of-the-art on two recently released benchmark datasets.
2 Related Work
Roadside monocular 3D detection helps smart transportation and smart city by enlarging the range of perception with a camera hung on a high infrastructure [56]. To foster research of roadside 3D detection, the community established large-scale datasets such as DAIR-V2X-I [61] and Rope3D [59]. As the roadside camera is typically fixed, multiple methods propose to exploit this prior knowledge to approach in roadside 3D detection [14, 40, 56, 58, 48]. For example, BEVHeight [56] trains a model to regress towards object height instead of object distance or depth, as done previously [18] which finds that regressing height is an easier task than regressing depth given the fixed camera pose. BEVSpread [48] further considers the position approximation error in the voxel pooling process based on BEVHeight. Different from existing methods, we identify a fruitful space to improve existing methods (Table 1) by exploiting a 2D detector (Fig. 1b) to produce 2D detections, using which to facilitate 3D detection (Fig. 1a). Further, we derive a scene prior by exploiting the fixed camera pose, and incorporate the scene prior to enhance 3D detection.
2D object detection aims to detect objects of predefined categories in a 2D image [10, 20]. Prevailing approaches train deep neural networks, which typically consist of a backbone to encode input images, and a detector-head for box regression and classification [36, 35, 22, 44]. Recently, using the transformer architecture significantly improves 2D detection [8, 65, 62]. In this work, we explore how to leverage a 2D detector [62, 44] to train better 3D detectors on roadside scenes. It is important to note that any 2D detectors can be used in our framework Pro3D.
Figure 2: We study three methods of prompt encoding and fusion, used to exploit a 2D detector to facilitate 3D detection. Design (a) concatenates feature maps extracted by the 2D detector and the 3D detector’s backbone. Design (b) extracts a feature vector based on a 2D detection’s coordinates, encodes it through a prompt encoder (implemented as an MLP), and attentively fuses it with the feature map extracted by the 3D detector’s backbone. Design (c) encodes a 2D detection, a 5-dim vector (coordinates $x$ and $y$, object width $w$ and height $h$, and the predicted class label) as the prompt, and attentively fuses the encoded detection with the feature map of the 3D detector’s backbone through the fusion module. Somewhat surprisingly, the third performs the best (Table 6)!
Prompting, enabled by large language models (LLMs), is first recognized in Natural Language Processing (NLP) [6, 39, 51]. Briefly, given an LLM, using a text prompt can instruct the LLM to produce desired text outputs. Inspired by this, the computer vision community endeavors to train large visual/multimodal models that can take text prompts to generate, restore and edit images [32, 23, 29]. Some works use images as prompts to accomplish vision tasks such as segmentation and inpainting [2, 49]. Beyond visual and text prompts, some recent works use more diverse information as prompts to guide models to produce desired output. For example, the Segment Anything Model (SAM) accepts bounding boxes, segmentation masks, and points as prompts towards interactive segmentation [15]. In this work, we propose the promptable 3D detector design (Pro3D), which trains a 3D detector that accepts prompts given by a pre-trained 2D detector. The prompt can be as simple as 2D detections, i.e., predicted box coorediates and class labels. Our Pro3D greatly boosts roadside monocular 3D detection performance (Fig. 1c, Table 1).
3 Promptable 3D Detector
We start with the problem definition of roadside monocular 3D object detection. Then, we introduce our Promptable 3D detector (Pro3D), including its crucial modules and a simple technique to obtain scene prior to improve detection performance. Lastly, we present insightful discussions.
Problem definition. Roadside monocular 3D detection requires detecting objects in a 2D RGB image and predicting their 3D information $(x,y,z,w,h,l,yaw)$, where $(x,y,z)$ is the object’s center location in the 3D world, $(w,h,l)$ denotes the width, height and length of the cuboid capturing the 3D object, and $yaw$ denotes the orientation of the cuboid on the BEV plane.
3.1 The Proposed Framework: Pro3D
Our Pro3D exploits a 2D detector for 3D detection and has two core modules to connect the two detectors (Fig. 1a): a prompt encoder and a fusion module. The former encodes 2D detections as new features to prompt the 3D detector. The latter attentively fuses the encoded prompts with the 3D detector backbone for 3D detection. The 3D detector head makes 3D predictions based on the fused features. We elaborate on the two modules below, followed by scene prior generation.
Prompt encoder treats the output of a 2D detector (either features or boxes) as prompt, and transforms it as a new feature which is exploited for 3D detections. The prompt encoder can be implemented as either a set of convolution layers or a multi-layer perceptron (MLP), depending on what are used as prompts and how to fuse (explained later in the fusion module). In our work, we compare three different prompts extracted from the 2D detector: (1) global feature maps represented as tensors, (2) local feature vectors at the spatial position specified by the predicted coordinates of 2D detections, and (3) the 2D detection consisting of box coordinates $x$ and $y$, width, height, and predicted class label. If feature maps are used as prompts, the prompt encoder can be a convolutional network that transforms the feature maps with appropriate dimensions (Fig. 2a), ensuring that the transformed feature maps can be fused with those of the 3D detector’s backbone. If a 2D detection is used as prompt, we represent the 2D box with a normalized positional encoding of its top-left and bottom-right corners, and its predicted class label as a class ID. We learn an MLP to transform them into a feature vector, using which to attentively fuse the feature map of the 3D detector backbone (details in supplement). Somewhat surprisingly, the third prompt performs the best (Table 6)! We explain next.
Remark 1: Using 2D detections as prompts yields better 3D detection performance than using features. We explain this by first noting different features extracted from a 2D detector: (1) global feature maps, and (2) local features at a specific position of the feature maps. Global feature maps contain all information from the input image, beneficial to object localization but lacking salient signals for semantic prediction. In contrast, local features may be informative for semantic prediction but may not contain spatial information for localization. Different from these features, the simplistic 2D detections (consisting of box coordinates, object sizes, and predicted labels) can well localize target objects in the input image and deliver their well-predicted labels. By explicitly incorporating the 2D detections, a 3D detector can be trained to lift them to the 3D space towards 3D detections. This simplifies the task of roadside monocular 3D detection and facilitates 3D detector training.
Figure 3: Diagram of our proposed fusion module, which attentively fuse the prompt $p$ (represented by its feature $F_{prompt}(p)$) and feature map $F(X)$ (from the 3D detector’s backbone) and ouptut the fused feature $f$.
Remark 2: Different from existing methods, our Pro3D explicitly leverages 2D detections. Leveraging 2D box annotations to facilitate 3D detector training is a de facto practice. One recent work similar to ours is MV2D [50], which trains a 3D detector from scratch for joint 2D and 3D detections. It is important to note that MV2D fuses RoI features extracted from the 2D detector with those of the 3D detector’s backbone for 3D detection. This is similar in spirit to our explored baseline method (Fig. 2b). Differently, our final method Pro3D explicitly leverages a 2D detector, without extracting its features, to facilitate 3D detection. Our experiments demonstrate that our Pro3D significantly outperforms MV2D (Table 1). Importantly, owing to the modular design, Pro3D can readily exploit any 2D detectors – using more advanced ones yields better 3D detection performance (Fig. 6).
Fusion module fuses the encoded prompt and the feature maps of the 3D detector backbone (Fig. 2). If the prompt is represented by a feature map extracted from the 2D detector, a straightforward fusion method is to concatenate these feature maps (Fig. 2a), as widely done in the literature [25, 31]. Yet, our final fusion module fuses feature vectors which represent 2D detections as prompts. To fuse the feature vectors, we take inspiration from the Transformer architecture [43, 15] and adopt an attention module. Refer to Fig. 3 for an illustration diagram. Concretely, for an input image $X$ which goes through the 3D detector backbone $F$, its feature map is denoted as $F(X)$. The prompt encoder $F_{prompt}$ transforms a prompt $p$ into a vector $F_{prompt}(p)$. The fusion model returns the final fused feature $f=\text{Norm}(\text{CrossAttn}(\text{query}=F(X),\text{key}=f_{2},\text{value}=f_{2}))$, where
[…]
$\text{query}=f_{1},\text{key}=F(X),\text{value}=F(X))))$
$f_{1}=\text{Norm}($ $\text{SelfAttn}(F_{\text{prompt}}(p)))$
Experiments demonstrate that such attentive fusion outperforms concatenation (Table 6).
Figure 4: Scene prior generation. We mask out objects of interest in training frames belonging to a specific scene, and average the remaining pixels across frames towards a “blank” background, which is our scene prior. A 3D detector that incorporates this scene prior (Fig. 1a) achieves remarkable improvements in roadside monocular 3D detection (Table 3).
Scene prior. Hung on a roadside perception infrastructure, the camera is fixed. Hence, the scene geometry in the captured video frames can be assumed to be stable. This inspires us to derive a scene prior from these frames. To generate a scene prior, we mask out objects (by running a 2D detector on the training frames belonging to a specific scene), and then average the remaining pixels across frames to obtain a “blank” background image for this scene, which is the scene prior (Fig. 4). In this work, we simply concatenate the scene prior with the RGB frame as a six-channel input to the 3D detector. We study the affects of using different scene priors generated in daytime and nighttime, respectively; results in Table 4 show that using either remarkably enhances performance.
Table 1: Benchmarking results on the DAIR-V2X-I dataset. Following [56], we report AP metrics at IoU=0.5, 0.25, 0.25 in BEV on the three superclasses (Vehicle, Cyclist, and Pedestrian) respectively. We mark the modalties (M) used in each method: C for RGB camera, L for LiDAR. As our Pro3D can take in any 3D detectors, we build it with BEVDepth, BEVHeight, and BEVSpread, respectively. Results show that Pro3D significantly improves these methods, e.g., it helps BEVHeight get $>$10 AP improvements on Pedestrian and Cyclist!
Method M Vehicle Cyclist Pedestrian
Easy Mid Hard Easy Mid Hard Easy Mid Hard
PointPillars [16] L 63.1 54.0 54.0 38.5 22.6 22.5 38.5 37.2 37.3
SECOND [54] L 71.5 54.0 54.0 54.7 31.1 31.2 55.2 52.5 52.5
MVXNet [42] L+C 71.0 53.7 53.8 54.1 30.8 31.1 55.8 54.5 54.4
FCOS3D [45] C 54.2 49.9 50.1 25.6 24.3 24.4 19.2 19.0 19.0
M3D-RPN [4] C 49.6 40.3 40.2 22.5 20.0 20.6 12.1 12.3 12.5
Kinematic3D [5] C 50.6 45.7 45.6 23.6 22.1 22.3 13.5 13.6 13.1
MonoDLE [28] C 52.3 48.2 48.1 24.1 23.5 23.6 14.6 14.4 14.5
MonoFlex [63] C 60.0 51.1 51.3 26.1 26.0 26.3 18.2 18.3 18.5
CaDDN [34] C 51.2 46.6 46.7 23.9 22.8 22.6 14.0 13.9 13.9
MonoNeRD [53] C 65.2 57.3 57.0 40.0 40.2 40.5 30.1 29.6 30.2
ImvoxelNet [37] C 44.8 37.6 37.6 21.1 13.6 13.2 6.8 6.7 6.7
BEVFormer [19] C 61.4 50.7 50.7 22.2 22.1 22.1 16.9 15.8 16.0
BEVFormer v2 [55] C 62.9 52.3 52.3 25.3 25.2 25.2 18.0 17.1 17.2
SparseBEV [21] C 65.6 55.1 52.6 30.0 30.3 30.0 18.4 17.9 17.5
MV2D [50] C 65.7 57.1 57.0 43.8 42.9 43.1 29.0 28.3 29.1
MonoGAE [57] C 84.6 75.9 74.2 44.0 47.6 46.8 25.7 24.3 24.4
BEVHeight++ [58] C 79.3 68.6 68.7 60.8 60.5 61.0 42.9 40.9 41.1
BEVDepth [18] C 75.5 63.6 63.7 55.7 55.5 55.3 35.0 33.4 33.3
Pro3D w/ BEVDepth C 79.6 71.8 71.7 65.8 65.8 65.3 46.0 44.7 44.9
performance gain C +4.1 +8.2 +8.0 +10.1 +10.3 +10.0 +11.0 +11.3 +11.6
BEVHeight [56] C 77.8 65.8 65.9 60.2 60.1 60.5 41.2 39.3 39.5
Pro3D w/ BEVHeight C 83.7 71.8 72.0 71.3 71.2 71.4 53.9 52.9 52.6
performance gain C +5.9 +6.0 +6.1 +11.1 +11.1 +10.9 +12.7 +13.6 +13.1
BEVSpread [48] C 79.1 66.8 66.9 62.6 63.5 63.8 46.5 44.5 44.7
Pro3D w/ BEVSpread C 85.2 73.1 73.6 73.1 73.0 73.1 55.1 54.3 54.1
performance gain C +6.1 +6.3 +6.7 +10.5 +9.5 +9.3 +8.6 +9.8 +9.4
3.2 Discussions on 2D Detector Training
Our Pro3D exploits a 2D detector (Fig. 1a). We can train this 2D detector using roadside 3D detection datasets. As such datasets contain only 3D cuboid annotations, we derive 2D box annotations by using the camera extrinsics and intrinsics to project the 3D cuboids onto the 2D image plane. In principle, our method can use any 2D detectors, e.g., pretraining a powerful “base” 2D detector and finetuning on the target dataset. Below we make three remarks on this matter.
Techniques of 2D detection are well established. 2D detection is more sufficiently explored than monocular 3D detection, as the former has been a fundamental problem in computer vision whereas the latter is relatively new. There are many excellent 2D detectors, such as FasterRCNN [36], YOLO [3, 44], DETR [8] and DINO [62]. For benchmarking, we use the recent transformer-based detector DINO [62]. We also test other 2D detectors in experiments.
More datasets of 2D detection are publicly available. As 2D detection has been actively studied since the past decade, many large-scale datasets have been released to the public. In contrast, for the relatively new problem of monocular 3D detection, there are less datasets which are also small in scale compared with the 2D counterpart. For example, the 2D detection dataset COCO [20] (published in 2014) annotates 330K images with 2D boxes for 80 classes, whereas the 3D detection dataset nuScenes [7] (published in 2020) “only” annotates 144K images with 3D cuboids for 23 classes. These publicly available 2D datasets allows pretraining stronger 2D detectors, using which our Pro3D yields better 3D detection performance (Fig. 6).
Joint training vs. stage-wise training. For 3D detection, most methods jointly train the 3D detector and a 2D detector by incorporating 2D annotations in an auxiliary loss [18, 24, 30, 41, 46, 45, 63, 11, 55, 50, 27]. Our Pro3D is a stage-wise training method, which turns out to outperform joint training (Table 7). We conjecture this is due to two reasons. First, 2D detector training is better explored with mature techniques than 3D detector training, so stage-wise training makes use of such techniques of 2D detector training. Second, joint training must strike a balance between various loss functions such as 3D cuboid regression, rotation regression and so on; yet, optimizing over all losses inevitably sacrifices 2D detection which is actually much easier than 3D detection. It is worth noting that stage-wise training allows easy leverage of external 2D datasets for pretraining a stronger 2D detector, exploiting which helps our Pro3D achieve even better performance (Fig. 6).
Figure 5: Visual comparison between the state-of-the-art method BEVHeight [56] and our Pro3D. Results show that Pro3D can (1) make better orientation predictions than BEVHeight (column-1), and (2) detect objects (missed by BEVHeight) that are too small in size (column-2), heavily occluded (column-3), and in the far field (column-4).
Table 2: Benchmarking results on the Rope3D dataset. Following [59], we use metrics AP and Rope at IoU=0.5 and 0.7. Built with detectors BEVHeight and BEVSpread, Pro3D achieves significant improvements. Notably, it improves the original BEVHeight and BEVSpread by $>$26 AP on Big Vehicle! Conclusions in Table 1 hold on this dataset.
Method IoU = 0.5 IoU = 0.7
Car Big Vehicle Car Big Vehicle
AP Rope AP Rope AP Rope AP Rope
M3D-RPN [4] 54.2 62.7 33.1 44.9 16.8 32.9 6.9 24.2
Kinematic3D [5] 50.6 58.9 37.6 48.1 17.7 32.9 6.1 22.9
MonoDLE [28] 51.7 60.4 40.3 50.1 17.7 32.9 6.1 22.9
MonoFlex [63] 60.3 66.9 37.3 48.0 33.8 36.1 10.1 26.2
BEVFormer [19] 50.6 58.8 34.6 45.2 24.6 38.7 10.1 25.6
SparseBEV [21] 54.7 62.3 39.8 48.4 29.9 41.2 14.6 30.0
BEVDepth [18] 69.6 74.7 45.0 54.6 42.6 53.1 21.5 35.8
BEVHeight++ [58] 76.1 80.9 50.1 59.9 47.0 57.7 24.4 39.6
BEVHeight [56] 74.6 78.7 48.9 57.7 45.7 55.6 23.1 37.0
Pro3D w/ BEVHeight 84.0 86.2 75.1 79.2 51.0 59.2 32.1 44.3
performance gain +9.4 +7.5 +26.2 +21.5 +5.3 +3.6 +11.0 +7.3
BEVSpread [48] 76.9 80.8 51.2 60.3 47.1 57.6 25.6 39.3
Pro3D w/ BEVSpread 86.3 87.6 77.6 81.0 53.1 60.6 33.9 45.6
performance gain +9.4 +6.8 +26.4 +20.7 +6.0 +3.0 +8.3 +6.3
4 Experiments
We validate our Pro3D and design choices through extensive experiments. First, we present experiment settings including datasets, evaluation metrics and implementations details. Then, we compare our Pro3D against prior works. Lastly, we conduct comprehensive ablation studies. We refer the reader to the supplement for more analyses.
4.1 Settings
Datasets. We use two publicly available datasets in experiments: DAIR-V2X-I [61], and Rope3D [59].
  • DAIR-V2X-I is the first large-scale dataset supporting the research of roadside monocular 3D detection. It provides 5K and 2K images for training and validation, respectively. Following [56], we report results on the validation set, with breakdown analysis w.r.t three superclasses, “vehicle”, “cyclist”, and “pedestrian”, at three difficulty levels (Easy, Medium, Hard).
  • Rope3D is another large-scale dataset that contains 50k images and annotations of $>$1.5M objects in diverse scenes. Following [56], we use its train-set (70%) to train models and the validation set (30%) for benchmarking.
Metrics. We introduce benchmarking metrics below.
[…]
  • Mean AP (mAP) [20] is used to report 2D detection performance on the image plane (Table 5 and Fig. 6). It first computes per-class average precision (AP) over IoU threshods from 0.5 to 0.95 with a step-size 0.05, and then returns the mean of per-class APs.
[…]
Compared methods. Our framework Pro3D can take in any 2D detectors and 3D detectors. To show the benefits of Pro3D, we build it with representative roadside 3D detectors, including BEVDepth [18], BEVHeight [56], and BEVSpread [48]. For benchmarking, we compare against more methods which are either published recently or representative in the literature 3D detection. We refer the reader to Table 1 and 2 for the compared methods.
Table 3: Incorporating scene prior remarkably improves roadside 3D detection. Interestingly, small objects of Cyclist and Pedestrian get more improvements than Vehicle.
Method Vehicle Cyclist Pedestrian
Easy Mid Hard Easy Mid Hard Easy Mid Hard
BEVHeight [56] 77.8 65.8 65.9 60.2 60.1 60.5 41.2 39.3 39.5
+ scene prior 79.2 67.7 67.8 62.5 62.3 62.7 44.3 42.5 42.7
performance gain +1.4 +1.9 +1.9 +2.3 +2.2 +2.2 +3.1 +3.2 +3.2
BEVSpread [48] 79.1 66.8 66.9 62.6 63.5 63.8 46.5 44.5 44.7
+ scene prior 80.1 68.5 68.6 64.5 65.4 65.6 49.6 47.5 47.7
performance gain +1.0 +1.7 +1.7 +1.9 +1.9 +1.8 +3.1 +3.0 +3.0
Pro3D w/ BEVSpread 84.3 71.5 72.0 71.4 71.3 71.5 52.2 51.6 51.4
+ scene prior 85.2 73.1 73.6 73.1 73.0 73.1 55.1 54.3 54.1
performance gain +0.9 +1.6 +1.6 +1.7 +1.7 +1.6 +2.9 +2.7 +2.7
Implementations. 3D detectors studied in our work mostly follow BEVHeight [56] and BEVSpread [48], which contain an image encoder with ResNet101 architecture [13], a BEV encoder with ResNet18 architecture, an FPN for proposal detection [54], and a detector head for the final output of 3D detections. It is worth noting that our Pro3D replaces the ResNet18-based BEV encoder with a single convolutional layer, removing the FPN as our Pro3D uses 2D detection already. This helps our Pro3D achieve faster inference than these previous mehtods (Table 9). Within our Pro3D framework, we report performance by using BEVDepth, BEVHeight and BEVSpread as the 3D detection module, respectively. As BEVHeight has the fastest inference speed (detailed in the supplement), we use it in our Pro3D for ablation study. For 2D detector, we train DINO [62] with the backbone ViT-S/16 as the 2D detector using the default hyperparameter setting. For all methods, the input image resolution is 864$\times$1536 and the initial voxel resolution is 1024$\times$1024$\times$1. We use random scaling and rotation augmentation during training, and train all models for 50 epochs (long enough for convergence) using AdamW optimizer [26] with learning rate 8e-4. We select the best checkpoint via validation. As a post-processing step, we remove detections that have confidence scores lower than 0.3. The supplementary material contains more detailed implementations, video demo, and code.
Table 4: Comparison of different scene priors for our Pro3D. Using scene prior generated by using either all training images or those captured at night enhances the final performance
Method Vehicle Cyclist Pedestrian
Easy Mid Hard Easy Mid Hard Easy Mid Hard
Pro3D w/ BEVSpread [48] 84.3 71.5 72.0 71.4 71.3 71.5 52.2 51.6 51.4
   +Scene priors (All) 85.2 73.1 73.6 73.1 73.0 73.1 55.1 54.3 54.1
   +Scene priors (Night) 85.0 72.7 73.0 72.7 72.5 72.6 53.9 53.6 53.4
4.2 Benchmarking Results
We study how much improvements our Pro3D brings to existing roadside 3D detectors on DAIR-V2X-I (Table 1) and Rope3D (Table 2). Fig. 5 visualizes results by Pro3D and BEVHeight (refer to the supplement for more visual comparisons). The results demonstrate that our Pro3D helps previous methods BEVDepth [18], BEVHeight [56], and BEVSpread [48] achieve significant performance gains. Moreover, Pro3D boosts 3D detection performance particularly on small objects of “cyclist” and “pedestrian”.
4.3 Ablation Study and Analysis
Incorporating scene prior improves 3D detection. We concatenate a testing image and the scene prior (a “blank” RGB background image as shonw in Fig. 4) as a 6-channel input to the 3D detector. Table 3 shows that adding the scene prior remarkably improves roadside 3D detection. To study how different scene prior affects the final performance, we generate scene priors over training frames captured at night. Table 4 shows that our method is quite robust to scene priors that have different illuminations.
Figure 6: We train different 2D detectors (YOLOV7 [44] and DINO [62]) on different amounts of 2D annotations (provided by nuImages [7]) for Pro3D. We measure the 3D detection by averaging AP over the three classes vehicle, cyclist and pedestrian. Training on more 2D annotations leads to better 2D detectors and importantly, increases 3D detection performance of Pro3D.
Exploiting a stronger 2D detection yields better 3D detection performance. We compare methods’ performance w.r.t 2D detection metrics in Table 5. This helps us understand that (1) 2D detection performance greatly reflects the 3D detection performance; (2) exploiting a stronger 2D detector helps our Pro3D achieve better 3D detection performance. We train stronger 2D detectors (e.g., YOLOV7 [44] and DINO [62]) on more 2D annotations from nuImages [7]. Fig. 6 shows 3D detection performance by our Pro3D versus 2D detection performance (by training different 2D detectors). Results show that exploiting stronger 2D detectors greatly improves 3D detection performance with our Pro3D.
Table 5: Comparison w.r.t 2D detection metric (mAP) on DAIR-V2X-I. For both 2D detector (DINO) and 3D detectors BEVDepth and BEVHeight, we train them on the training set of DARI-V2X-I. To evaluate 3D detectors w.r.t 2D metrics, we project their 3D detections onto the 2D image plane to derive their predicted 2D boxes. Compared to BEVDepth and BEVHeight, DINO achieves $>$10 mAP higher on all classes and $\sim$20 mAP higher on Pedestrian! Owing to the effective leverage of the 2D detector DINO, our Pro3D significantly boosts performance of the compared methods.
Method Vehicle Cyclist Pedestrian
Easy Mid Hard Easy Mid Hard Easy Mid Hard
2D detector (DINO) [62] 74.3 75.6 75.6 57.3 54.7 54.9 50.6 51.6 51.6
BEVDepth [18] 62.9 64.8 64.2 41.5 44.8 44.9 32.9 33.1 33.0
Pro3D w/ BEVDepth 68.0 68.3 68.9 45.9 47.7 47.6 42.1 42.6 42.6
BEVHeight [56] 63.5 64.7 64.5 41.4 44.6 44.7 33.6 33.7 33.9
Pro3D w/ BEVHeight 68.1 68.8 68.8 47.9 49.5 49.4 42.8 43.5 43.5
BEVSpread [48] 64.1 65.2 64.6 41.7 45.1 44.9 33.9 34.1 34.3
Pro3D w/ BEVSpread 68.7 69.4 69.2 48.7 50.3 50.3 44.1 44.9 44.9
Table 6: Comparison of different prompting methods. As a baseline, we train BEVHeight with scene prior (Table 3). First, all prompting methods achieve remarkable improvements over BEVHeight. Moreover, using the center of predicted 2D box coordinates (“fuse 2D centers”) as prompts outperforms feature prompts. Lastly, additionally incorporating predicted labels brings further improvements.
Method Vehicle Cyclist Pedestrian
Easy Mid Hard Easy Mid Hard Easy Mid Hard
BEVHeight [56] 79.2 67.7 67.8 62.5 62.3 62.7 44.3 42.5 42.7
+ concatenate feaMaps 79.6 68.0 67.7 63.0 62.6 62.7 44.2 42.6 42.8
+ attentively fuse local vectors 79.6 68.1 68.1 64.0 63.8 63.9 44.1 43.3 43.4
+ attentively fuse box centers 79.8 68.3 68.3 65.2 65.0 65.1 46.8 45.7 45.8
+ predicted labels 80.2 68.6 68.6 66.1 65.5 65.6 47.5 46.4 46.5
+ attentively fuse 2D boxes 80.1 68.6 68.5 66.7 66.1 66.2 48.2 47.1 47.2
+ predicted labels 80.5 68.9 68.9 68.9 68.0 68.2 50.7 50.0 50.1
performance gain +1.3 +1.2 +1.1 +6.4 +5.7 +5.5 +6.4 +7.5 +7.4
What to use as prompts. To exploit the 2D detector to help train 3D detectors, we study what information from it can be used as prompts, including its intermediate feature map, feature vector located at a 2D detection, and even the simplistic 2D detection output (i.e., predicted box coordinates and class label). Note that they require different fusion methods to accommodate their different dimensions. For example, when using feature maps, one can concatenate them with feature of the encoder in the 3D detector, while using feature vectors or detections output require fusing vectors with the feature maps. Therefore, we use concatenation to fuse feature maps, and an attention module to fuse the feature vectors. Table 6 lists the results. Surprisingly, using the simple detections (2D boxes and class labels) as prompts performs better than feature maps and feature vectors located at 2D detections, where 2D boxes as prompts achieves [68.5, 66.2, 47.2] AP for Hard examples of the three classes [vehicle, cyclist, pedestrian]. Furthermore, additionally incorporating the predicted class labels in prompts performs the best, boosting to [68.9, 68.2, 50.1] AP on Hard examples of the three classes! In hindsight, we see that fusing box coordinates and labels provides both semantic and spatial information for the 3D detector to leverage, facilitating 3D detection. In contrast, fusing feature maps does not pinpoint the location of the objects and hence does not simplify 3D detector training.
Table 7: Comparison of stage-wise training (SW) vs. joint training (JT) for our Pro3D. We report AP metrics at IoU=0.5, 0.25, 0.25 in BEV. As a baseline, we train Pro3D with BEVHeight and scene prior. It is worth noting that stage-wise training allows us to train separate 2D and 3D detector backbones, i.e., their backbones are not shared (NS). Due to this, we can exploit more 2D annotated data (MD) to train better 2D detectors, e.g., using nuImages [7] to pretrain a 2D detector. Results convincingly show that stage-wise training outperforms joint training for roadside monocular 3D detection. With NS and MD, our Pro3D yields more notable improvements.
Method Vehicle Cyclist Pedestrian
Easy Mid Hard Easy Mid Hard Easy Mid Hard
Pro3D 80.5 68.9 68.9 68.9 68.0 68.2 50.7 50.0 50.1
+ JT 80.6 68.8 68.9 68.9 68.2 68.3 50.6 50.1 50.3
+ SW 81.5 69.3 69.6 69.5 69.1 69.2 51.3 50.7 50.3
+ SW, NS 82.3 69.7 70.0 70.0 69.9 70.1 52.1 51.0 50.8
+ SW, NS, MD 82.7 70.1 70.3 70.5 70.4 70.6 52.8 51.7 51.3
perf. gain +2.2 +1.2 +1.4 +1.6 +2.4 +2.4 +2.1 +1.7 +1.2
Stage-wise training vs. joint training. We test different training paradigms in our Pro3D, e.g., (1) joint end-to-end training the 3D detector using both 2D and 3D losses, and (2) stage-wise training the 2D detector first and then the 3D detector. Note that stage-wise training allows us to train separate 2D and 3D detector backbones, i.e., their backbones are not shared (NS). This further makes it convenient to exploit more 2D annotated data (MD) to train better 2D detectors, e.g., using nuImages [7] to pretrain a 2D detector. Results in Table 7 show that stage-wise training outperforms joint training for roadside monocular 3D detection.
How to train 3D detectors with 2D detection prompts. The Pro3D uses 2D detection to prompt the 3D detector to make 3D predictions. To train the 3D detector, we can either use ground-truth 2D boxes or predicted ones as prompts. We compare their performance w.r.t AP in BEV on the DAIR-V2X-I benchmark. Results in Table 8 show that directly using the predicted 2D boxes (with predicted class labels) as prompts performs better.
Table 8: How to train Pro3D. When training the 3D detector, we can construct prompts as either ground-truth 2D boxes (grnd) or predicted ones (pred). We compare them on DAIR-V2X-I w.r.t AP in BEV. Results show that using latter performs better.
Method Vehicle Cyclist Pedestrian
Easy Mid Hard Easy Mid Hard Easy Mid Hard
Pro3D w/ grnd 82.7 70.1 70.3 70.5 70.4 70.6 52.8 51.7 51.3
Pro3D w/ pred 82.9 70.7 70.8 71.3 71.2 71.4 53.9 52.9 52.6
Latency and parameters. Recall that our Pro3D achieves the best performance by using separate 2D and 3D detectors without sharing the visual encoder (Table 7). We study this model in Table 9 w.r.t model parameters, the wall-clock time, and FPS in processing a single RGB frame. Here, we run the 2D detector (either YOLOV7 or DINO) and the 3D detector encoder in parallel. Note that our Pro3D does not require an FPN for proposal detection because the 2D detector provides pinpointed detections already. Our Pro3D also replaces ResNet18 encoder as adopted in prior methods [56] with a single convolutional layer (Sec. 4.1). These changes make our Pro3D faster than the original BEVHeight in inference, regardless exploiting either YOLOV7 [44] or DINO [56] as the 2D detector. In summary, our Pro3D model “w/ YOLOV7” boosts 3D detection performance by 7.9 AP over the prior art BEVHeight, with faster inference speed (15.4 vs. 13.0 FPS) and fewer parameters (86.6M vs. 94.7M)!
Table 9: Comparison of inference time and parameters between our Pro3D and BEVHeight. AP means the average AP in BEV of all the three classes in DAIR-V2X-I; “#param” counts the parameters (in million); time is the wall-clock in ms, FPS is frames per second. We use two different 2D detectors in our Pro3D, namely YOLOV7 and DINO. We list the number of parameters of 3D detector plus 2D detector. Our model Pro3D “w/ YOLOV7” boosts 3D detection performance by 7.9 AP over the prior art BEVHeight, with faster inference speed (15.4 vs. 13.0 FPS) and fewer parameters (86.6M vs. 94.7M)!
Methods AP #param. time FPS
BEVHeight [56] 57.1 94.7 77 13.0
Pro3D w/ BEVHeight + YOLOV7 65.0 82.1+4.5 65 15.4
Pro3D w/ BEVHeight + DINO 66.9 82.1+21.2 73 13.7
4.4 Societal Impacts and Limitations
Societal impacts. While our method achieves significantly better performance than prior works on the benchmark datasets of roadside 3D object detection, we note potential risks if directly using our method for vehicle-vehicle communication and vehicle-infrastructure cooperative perception. Real-world applications should treat our approach as a module and evaluate the whole system for ultimate goals. Methods developed with existing datasets might be biased towards specific cities and countries. For example, in some countries, vehicles are running on the different side of the road from what are presented in the datasets. These are potentially negative societal impacts.
Limitations and future work. Our Pro3D exploits a scene prior; if camera pose is changed (e.g., due to camera replacement or strong wind), using outdated scene prior might negatively affect performance. Moreover, Pro3D outputs a 3D detection given a 2D detection. As 2D detections can be false positives, doing so inevitably hallucinates 3D detections which certainly are incorrect. Future work will analyze this aspect in more depth.
5 Conclusion
We address the problem of roadside monocular 3D detection and propose a novel design, Promptable 3D detector (Pro3D), which leverages a pre-trained 2D detector to facilitate 3D detector training. Our Pro3D runs a 2D detector to obtain 2D detections, and uses them to prompt the 3D detector to make 3D predictions. We present novel modules to encode 2D detections and fuse them with features from the 3D detector’s backbone. Moreover, we present a simple method to obtain a scene prior, by incorporating which our Pro3D achieves further improvements. We validate Pro3D through extensive experiments and ablation studies. Comprehensive results show that Pro3D significantly outperforms prior works on publicly available benchmarks.
Supplementary Material
Outline
This document supplements the main paper with further details of Pro3D and more visualizations. Below is the outline of this document.
  • Section A. We provide implementation details of the prompt encoder in Pro3D.
  • Section B. We discuss our loss design in Pro3D.
  • Section C. We evaluate models’ performance under occlusion and truncation scenarios.
  • Section D. We study the robustness of Pro3D to camera noises.
  • Section E. We compare the inference speed of Pro3D using different 3D detectors.
  • Section F. We provide our code.
  • Section G. We provide more visual results, including a video demo.
Appendix A Details of Prompt Encoder
The prompt encoder transforms the 2D detection outputs to features through three steps: (1) normalize the top-left and bottom-right coordinates of a 2D detection box, denoted as $A\in{\mathbb{R}}^{2\times 2}$; (2) multiply $A$ by a random matrix $B\in{\mathbb{R}}^{2\times 512}$ whose entries are initialized using a Gaussian distribution, and then add a learnable matrix $C\in{\mathbb{R}}^{2\times 512}$ to get $D=A\times B+C$; (3) repeat the predicted class label (represented as a unique class ID) 512 times to get a 512-dimensional vector, and concatenate it with $D$ to obtain the final feature prompts $E\in{\mathbb{R}}^{3\times 512}$.
Figure 7: Comparison of two grouping methods in the loss design: (a) appearance-based grouping, and (b) functionality-based grouping. The former is commonly adopted in existing roadside 3D detectors that group classes according to their appearance similarity and use multi-head for prediction (each head is responsible for $K$-way fine-grained prediction. The latter is our proposed method that groups classes according to their functionalities and uses multi-head for prediction – each head is a single super-class predictor, which merges {car, van, truck, bus} into vehicle, and {bicyclist, tricyclist, motorcyclist, barrowlist} into cyclist.
Table 10: Results of occluded and truncated obstacles. We split occlusions and truncations with different ranges (0$\sim$50and 50$\sim$100%), respectively. Our approach Pro3D resoundingly outperforms BEVHeight [56] and BEVSpread [48] on all the three superclasses in each range.
(a) Occlusion of 0$\sim$50%
Methods Vehicle Cyclist Pedestrian
BEVHeight 75.6 66.8 48.9
Pro3D w/ BEVHeight 80.8 76.6 58.0
BEVSpread 77.5 67.7 51.2
Pro3D w/ BEVSpread 83.2 76.8 60.1
(b) Occlusion of 50$\sim$100%
Methods Vehicle Cyclist Pedestrian
BEVHeight 51.2 36.9 19.2
Pro3D w/ BEVHeight 58.0 42.3 28.1
BEVSpread 53.3 38.6 19.9
Pro3D w/ BEVSpread 60.1 44.6 28.5
(c) Truncation of 0$\sim$50%
Methods Vehicle Cyclist Pedestrian
BEVHeight 76.2 68.0 50.3
Pro3D w/ BEVHeight 83.9 76.8 60.0
BEVSpread 78.8 69.3 52.6
Pro3D w/ BEVSpread 85.0 77.9 61.6
(d) Truncation of 50$\sim$100%
Methods Vehicle Cyclist Pedestrian
BEVHeight 38.7 18.6 10.8
Pro3D w/ BEVHeight 45.5 27.3 20.5
BEVSpread 38.9 19.1 11.2
Pro3D w/ BEVSpread 45.6 27.6 20.6
Table 11: Ablation on class-grouping strategies (w.r.t metric AP in BEV). Grouping classes differently with multi-head design (each detector head makes predictions within a specific superclass) makes a difference in training and yields detectors performing differently. Briefly, instead of appearance-based grouping (AG) which groups classes w.r.t object appearance as done by BEVHeight, we do functionality-based grouping (FG), e.g., merging {car, van, truck, bus} into the superclass vehicle. To justify the superiority of FG over AG, we replace the original AG of BEVHeight [56] and BEVSpread [48] with FG, particularly improving on the vehicle class (cf. “BEVHeight w/ FG” vs. “BEVHeight w/ AG”, “BEVSpread w/ FG” vs. “BEVSpread w/ AG”).
Method Vehicle Cyclist Pedestrian
Easy Mid Hard Easy Mid Hard Easy Mid Hard
BEVHeight w/ AG 77.8 65.8 65.9 60.2 60.1 60.5 41.2 39.3 39.5
BEVHeight w/ FG 80.0 67.9 68.1 60.5 60.3 60.7 41.4 39.4 39.7
performance gain +2.2 +2.1 +2.2 +0.3 +0.2 +0.2 +0.2 +0.1 +0.2
BEVSpread w/ AG 79.1 66.8 66.9 62.6 63.5 63.8 46.5 44.5 44.7
BEVSpread w/ FG 81.1 68.8 69.0 62.7 63.7 63.9 46.6 44.7 44.8
performance gain +2.0 +2.0 +2.1 +0.1 +0.2 +0.1 +0.1 +0.2 +0.1
Appendix B Loss Design
Class grouping is a common strategy in 3D object detector training. It merges classes into some superclasses, allowing trained features to be shared within superclasses [64]. Moreover, multi-head design follows along with groups of classes for discriminative predictions across fine-grained classes [60]. BEVHeight [56] and BEVSpread [48] adopt class grouping and multi-head design by considering the similarities among object appearance, e.g., size and shape in the 3D world. For example, on DAIR-V2X-I, they create three superclasses that merge {truck, bus}, {car, van}, and {bicyclist, tricyclist, motorcyclist, barrowlist}, respectively, along with the origin pedestrian class. The rationale of such grouping is based on object appearance, e.g., both truck and bus have large size. We call this grouping strategy appearance-based grouping. With such a group strategy, they learn a four-head detector with each making predictions for the corresponding superclass. Differently, we adopt functionality-based grouping, which merges {car, van, truck, bus} into vehicle, and {bicyclist, tricyclist, motorcyclist, barrowlist} into cyclist. This is motivated by the fact that vehicles, cyclists and pedestrians appear at relatively fixed regions on the roadside images. We expect this grouping strategy, or using the location prior of different classes, to facilitate training 3D object detectors on roadside images. Fig. 7 compares the two grouping methods. Perhaps from the same functionality perspective, the DAIR-V2X-I benchmark also merges classes into such superclasses for evaluation. For fair comparison, we use our functionality-based grouping loss to train the 3D detector (e.g., BEVHeight or BEVSpread) within our Pro3D framework to compare with the original 3D detectors. Our extensive study shows that trainig with the functionality-based grouping loss remarkably improves detection performance (Table 11).
Appendix C Analysis of Occlusion and Truncation
Objects might be occluded by others or truncated around the image borders. We carry out breakdown analysis how detectors perform in these scenarios. The DAIR-V2X-I dataset provides occlusion and truncation tags. We split occlusions and truncations with 0$\sim$50and 50$\sim$100for different superclasses on [61], respectively. As shown in Table 10, our method gets better APs on the all superclasses in different ranges of occlusions and truncations. This shows that our method Pro3D is more robust than BEVHeight [56] and BEVSpread [48].
Appendix D Robustness to Noises of Camera Pose
We consider the robustness of Pro3D to noises in camera calibration parameters. We follow BEVHeight [56] to add noises on camera pitch and roll. Table 12 compares the robustness of BEVHeight [56], BEVSpread [48] and our Pro3D. Clearly, Pro3D is more robust than prior works to camera noises.
Table 12: Robustness to camera noise. We follow BEVHeight [56] to simulate calibration errors by adding noises to camera pitch and roll. Results show that Pro3D is more robust than BEVHeight and BEVSPread [48] to these camera pose noises.
Method Vehicle Cyclist Pedestrian
Easy Mid Hard Easy Mid Hard Easy Mid Hard
BEVHeight 77.8 65.8 65.9 60.2 60.1 60.5 41.2 39.3 39.5
+ noise 62.5 49.9 50.7 45.4 43.4 43.8 34.5 32.6 32.7
performance degradation -15.3 -15.9 -15.2 -14.8 -16.7 -16.7 -6.7 -6.7 -6.8
Pro3D w/ BEVHeight 83.7 71.8 72.0 71.3 71.2 71.4 53.9 52.9 52.6
+ noise 72.5 59.9 60.3 59.0 60.0 61.9 47.9 46.8 46.5
performance degradation -11.2 -11.9 -11.7 -12.3 -11.2 -9.5 -6.0 -6.1 -6.1
BEVSpread 79.1 66.8 66.9 62.6 63.5 63.8 46.5 44.5 44.7
+ noise 64.2 51.6 51.4 48.4 48.8 49.4 40.3 37.6 38.6
performance degradation -14.9 -15.2 -15.5 -14.2 -14.7 -14.4 -6.2 -6.9 -6.1
Pro3D w/ BEVSpread 85.2 73.1 73.6 73.1 73.0 73.1 55.1 54.3 54.1
+ noise 73.1 60.3 61.7 60.4 61.8 63.1 48.9 48.0 47.8
performance degradation -12.1 -12.8 -11.9 -12.7 -11.2 -10.0 -6.2 -6.3 -6.3
Table 13: Comparison of inference time (ms) between different methods. BEVHeight [56] has the fastest inference time comparing with other three methods. Here, we build our Pro3D using it as the 3D detector. It is worth noting that we adopt the functionality-based grouping loss, which induces less parameters compared to the wide-used appearance-based grouping loss (Section B). This makes our Pro3D obtain the fastest inference speed among these methods.
Method BEVDepth BEVHeight BEVHeight++ BEVSpread Pro3D
time(ms) 82 77 92 90 65
Appendix E Comparison on the Inference Speed
We compare the inference time of different methods in Table 13. BEVHeight [56] gets the fastest inference time in previous work, so we use it in our ablation studies. Moreover, our method Pro3D outperforms other works in inference speed, where we use YOLOV7 as the 2D detector in Pro3D.
Appendix F Open-Source Code
Code. We include our self-contained codebase (refer to the zip file Pro3D-main) as a part of the supplementary material. Please refer to README.md for instructions how to use the code. We do not include model weights in the supplementary material as they are too larger than the space limit (100MB). We will open source our code and release our trained models to foster research.
License. We release open-source code under the MIT License to foster future research in this field.
Requirement. Running our code requires some common packages. We installed Python and most packages through Anaconda. A few other packages might not be installed automatically, such as Pandas, torchvision, and PyTorch, which are required to run our code. Below are the versions of Python and PyTorch used in our work:
  • Python version: 3.8.0 [GCC 7.5.0]
  • PyTorch version: 1.8.1
Appendix G More Visualizations and Video Demo
Figure 8 visualizes more results on the DAIR-V2X-I dataset [61]. Pro3D not only has better orientation predictions for vehicles but also can detect hard objects such as infrequently-seen ambulances, occluded pedestrians and cyclists.
Video demo. We include a video demo as a part of the supplementary material, named video-demo.mp4. In this video demo, we compare our Pro3D with the state-of-the-art method BEVSpread [48]. The demo clearly shows that our Pro3D performs much better than BEVSpread in roadside monocular 3D detection.
Figure 8: More visualizations between the state-of-the-art method BEVHeight [56] and our Pro3D.